# Cospectral vertices, Walk-regular planar graphs and the echolocation problem

### v1

@@ Abstract.
We study cospectral vertices on finite graphs in relation to the echolocation problem on Riemannian manifolds. First, We prove a computationally simple criterion to determine whether two vertices are cospectral. Then, we use this criterion in conjunction with a computer search to find minimal examples of various types of graphs on which cospectral but non-similar vertices exist, including minimal walk-regular non-vertex-transitive graphs, which turn out to be non-planar. Moreover, as our main result, we classify all finite walk-regular planar graphs, proving that such graphs must be vertex-transitive.
@@ 1. Introduction
The study of cospectral vertices has gained more attention recently, (see e.g., [God12, GS17, ${}^{+}$19, EKL19, KSSW20]), especially due to their significant role in quantum information theory. It has been shown [God12] that for a perfect quantum state transfer to occur between two vertices in an interacting qubits network, it is necessary for these vertices to be cospectral.
@@
Smooth analogs of walk-regular graphs were studied by Wang and the last two authors in [WWX23]. There we show that every compact two-dimensional surface satisfying the smooth version of walk-regularity is also homogeneous—the isometry group acts transitively on the points in the surface. The assumption that our manifold is two-dimensional is essential to the proof, and it is yet unknown if there are any higher-dimensional manifolds that are walk-regular but not homogeneous. In addition, none of the walk-regular graphs we found are planar. This leads us to consider planar walk-regular graphs.
@@
Every walk-regular planar graph with degree $d\geq 3$ is isomorphic to the polyhedral graphs of one of the following:
@@
A direct consequence of the classical theorem of Mani [Man71] says that all planar vertex-transitive finite graphs with degree at least 3 are exactly those listed above. Herbert and Wilfried [FI79] employed modified Schläfli symbol to give a direct proof of this fact. Theorem 1.5 is a refinement of this result of Mani and Herbert–Wilfried, as walk-regularity is, in general, strictly weaker than vertex-transitivity. The modified Schläfli symbol again plays a crucial role in our proof of Theorem 1.5.
Figure 1. All the 13 semi-regular polyhedrons. The graphics are due to Wikipedia user Cyp [Cyp12].
@@
In Section 2, we prove Proposition 1.1. In Section 3, we give a short overview of our computer search and present a few of our minimal examples. The code in full is available on GitHub along with a repository of minimal examples (https://github.com/ewy-man/graph-echolocation). Section 4 is dedicated to the proof of Theorem 1.5.
@@ Acknowledgements
Xi was supported by the National Key Research and Development Program of China No. 2022YFA1007200 and NSF China Grant No. 12171424. Kong was supported by NSF China Grant No. 12301109 and the Fundamental Research Funds for the Central Universities of China No. YJ202343. Wyman was supported by NSF grant DMS-2204397. The second listed author extends a thanks to Richard Lange for his very patient guidance through the process of setting up a GitHub repository.
@@ 2. Closed walks of bounded length determine if two vertices are cospectral
We note that the number $n-1$ of counts is sharp, since the adjacency matrix $A$ of a simple path with $n$ vertices for every odd $n=2\ell+1$ (i.e., $A_{i,j}=\delta_{|i-j|=1}$) as well as the adjacency matrix $A$ of a simple path with $n$ vertices and a self-loop attached to the last vertex for every even $n=2\ell$ (i.e., $A_{i,j}=\delta_{|i-j|=1}+\delta_{i=j=n}$) satisfies
@@ 3. Minimal examples
The code is written in the Julia language (see [BEKS17, ${}^{+}$21]) and contains a number of tools:
@@
    A number of other convenience functions, such as those that determine if a graph is walk-regular, or if there exist non-similar cospectral pairs.
@@ 3.3. Minimal walk-regular, non-vertex-transitive graphs
There are no walk-regular, non-vertex-transitive graphs which have less than 12 vertices. There are exactly four minimal examples with twelve vertices, one each of degree 4, 5, 6, and 7. Again, these minimal examples are connected, and so can be put into edge-complement pairs; the examples of degrees $4$ and $7$ are edge complements, and so are the examples of $5$ and $6$. The degree $4$ and $5$ examples are depicted in Figure 3.
@@ 4. Walk-regular planar graphs
Here we will set up the tools we will need in order to classify the planar walk-regular graphs and prove Theorem 1.5.
@@
Using Euler’s formula $n+f-e=2$ and $e=nd/2$, we get
(4.1) $\frac{1}{n}\sum_{v}p(v)=\frac{d-2}{2}+\frac{2}{n}$
We let $k(v)=\#\{1\leq i\leq d:r_{i}(v)=r_{1}(v)\}$. An easy observation is
@@
In a walk-regular planar graph with $d\geq 3$,
[…]
Proof.
First we let $r_{1}=\min_{v}r_{1}(v)$. Then it follows from (4.1) that
$\frac{d}{r_{1}}\geq\frac{d-2}{2}+\frac{2}{n}.$
Thus $r_{1}\in\{3,4,5\}$ for $d=3$, and $r_{1}=3$ for $d=4$ or $5$. Let $v_{1}$ be a vertex such that $r_{1}(v_{1})=r_{1}$. Suppose that there is a vertex $v_{2}$ that satisfies $r_{1}(v_{2})>r_{1}$. If $r_{1}\in\{3,5\}$, we have
$A^{r_{1}}_{v_{2},v_{2}}=0<2k(v_{1})=A^{r_{1}}_{v_{1},v_{1}}.$
If $r_{1}=4$, then $d=3$ and
$A^{4}_{v_{2},v_{2}}=15<15+2k(v_{1})=A^{4}_{v_{1},v_{1}}.$
In any case, $v_{1}$ and $v_{2}$ are not cospectral. Hence $r_{1}(v)\equiv r_{1}$ provided that the graph is walk-regular. By a similar closed walk counting argument, we conclude that
$A_{v,v}^{r_{1}}-A_{v^{\prime},v^{\prime}}^{r_{1}}=2k(v)-2k(v^{\prime})$
for any pair of vertices $v,v^{\prime}$, and thus $k(v)\equiv k$ by walk-regularity. ∎
In order to prove Theorem 1.5, we shall prove the following lemma.
Lemma 4.2.
For any walk-regular planar graph with degree $d\geq 3$, the modified Schläfli symbol is constant over all vertices, that is
@@
However, the proof of this lemma appears to be quite delicate. The 3-regular planar graph, as illustrated in Figure 5, features cospectral vertices $v_{1}$ and $v_{2}$ with $r(v_{1})=(3,5,5)$ but $r(v_{2})=(3,5,6)$. Despite vertex $v_{1}$ having two neighboring pentagons and $v_{2}$ only one, the adjacent triangle $\{v_{5},v_{7},v_{9}\}$ contributes two additional closed walks of length 5 to $v_{2}$. Moreover, this graph is ‘almost’ walk-regular, with all vertices except $v_{4}$ being cospectral. Therefore, establishing (4.2) succinctly appears challenging. Instead, we justify (4.2) in each individual case.
We remark that, once we prove Lemma 4.2, it is routine to check that Theorem 1.5 holds in each case. Indeed, the numbers of vertices, edges, and faces are determined by Lemma 4.2 and Euler’s identity. The graph can then be reconstructed in a deterministic fashion, starting with an arbitrary polygon as the outer polygon. We illustrate this fact for the most complicated case when $r(v)\equiv(3,4,5,4)$ in Figure 4. The only exception is the case when $r(v)\equiv(3,4,4,4)$, for which one runs into two possible reconstructions, the polyhedral graphs of the $(3,4,4,4)$-solid and twisted $(3,4,4,4)$-solid. However, the twisted one is not walk-regular. See Case (13) in the proof.
@@
Figure 5. A 3-regular planar graph on 10 vertices.
Let $v_{1}$ be a vertex such that $p(v_{1})=\max_{v}p(v)$. In view of Lemma 4.1, to prove Lemma 4.2, we only need to discuss the sixteen cases enumerated below. For the convenience of the reader, we have also included an exhaustive list of the possible walk-regular planar graphs belonging to each case. From now on and throughout the paper, for the purpose of simplicity, we shall not distinguish between a solid and its polyhedral graph.
@@
(16) 5, 3, $\leq$ 3
Of these cases, (2) and (8) are simply impossible to realize. It is also impossible to realize case (16), since
$p(v)\leq\frac{1}{3}+\frac{1}{3}+\frac{1}{3}+\frac{1}{4}+\frac{1}{4}=\frac{3}{2}$
violates (4.1). One quickly sees the symbols $r(v)$ of the graphs realizing cases (1), (4), (7), (10), and (14) must be constant.
Next, we focus on cases (5), (11), (9), and (15), which are easier to resolve. The remaining cases (3), (6), (12), and (13) are more involved. We will treat these cases after.
@@
($d=3,r_{1}=4,k=2$) For each $m=r_{3}(v_{1})\geq 5$, the face $f_{3}(v_{1})$ surrounded by $m$ quadrangular faces determines the unique pattern up to isomorphism. In this case, $r(v)\equiv(4,4,m)$ and the graph is isomorphic to the $m$-gonal prism.
@@
($d=4,r_{1}=3,k=3$) For each $m=r_{4}(v_{1})\geq 4$, the face $f_{4}(v_{1})$ surrounded by $m$ triangular faces determines the unique pattern up to isomorphism. In this case, $r(v)\equiv(3,3,3,m)$ and the graph is isomorphic to the $m$-gonal antiprism.
Starting from the next case, we often need to count the closed walks of length $m$ for some $m\geq 4$ and two adjacent vertices $v,v^{\prime}$ with $r(v)=r(v_{1})\neq r(v^{\prime})$. Our strategy of counting is to consider only the $\lfloor\frac{m}{2}\rfloor$-neighborhood of $\{v,v^{\prime}\}$, which turns out to be determined by the values of $d,r_{1},k$ and $r(v_{1})$, with at most two possibilities.

### v2

@@ Abstract.
We study cospectral vertices on finite graphs in relation to the echolocation problem on Riemannian manifolds. First, We prove a computationally simple criterion to determine whether two vertices are cospectral. Then, we use this criterion in conjunction with a computer search to find minimal examples of various types of graphs on which cospectral but non-similar vertices exist, including minimal walk-regular non-vertex-transitive graphs, which turn out to be non-planar. Moreover, as our main result, we classify all finite 3-connected walk-regular planar graphs, proving that such graphs must be vertex-transitive.
@@ 1. Introduction
The study of cospectral vertices has gained more attention recently, (see e.g., [God12, GS17, CCT+19, EKL19, KSSW20]), especially due to their significant role in quantum information theory. It has been shown [God12] that for a perfect quantum state transfer to occur between two vertices in an interacting qubits network, it is necessary for these vertices to be cospectral.
@@
Smooth analogs of walk-regular graphs were studied by Wang and the last two authors in [WWX23]. There they show that every compact two-dimensional surface satisfying the smooth version of walk-regularity is also homogeneous—the isometry group acts transitively on the points in the surface. The assumption that our manifold is two-dimensional is essential to the proof, and it is yet unknown if there are any higher-dimensional manifolds that are walk-regular but not homogeneous. In addition, none of the walk-regular graphs we found are planar. This leads us to consider planar walk-regular graphs.
@@
Every 3-connected walk-regular planar graph is isomorphic to the polyhedral graphs of one of the following:
@@
Figure 1. All the 13 semi-regular polyhedrons. The graphics are due to Wikipedia user Cyp [Cyp12], used with permission.
A direct consequence of the classical theorem of Mani [Man71] says that all 3-connected planar vertex-transitive finite graphs are exactly those listed above. Fleischner and Imrich [FI79] employed the modified Schläfli symbol to give a direct proof of this fact. Theorem 1.5 is a refinement of this result of Mani and Fleischner–Imrich, as walk-regularity is, in general, strictly weaker than vertex-transitivity. The modified Schläfli symbol again plays a crucial role in our proof of Theorem 1.5. To be more detailed, we first attribute the counting of short closed walks to the counting of short cycles, and by some delicate arguments, we prove that in most 3-connected walk-regular planar graphs, every short cycle without chords must be a face, further demonstrating that the modified Schläfli symbol at each vertex is identical.
Remark 1.6.
Although the planarity assumption might seem excessively restrictive, it is sharp in the sense that there exists a walk-regular toroidal graph that is not vertex-transitive. See Figure 4.
@@
In Section 2, we prove Proposition 1.1. In Section 3, we give a short overview of our computer search and present a few of our minimal examples. The code is available in full on GitHub, along with a repository of minimal examples (https://github.com/ewy-man/graph-echolocation). Section 4 is dedicated to the proof of Theorem 1.5.
@@ Acknowledgements
Xi was supported by the National Key Research and Development Program of China No. 2022YFA1007200 and NSF China Grant No. 12171424. Kong was supported by NSF China Grant No. 12301109 and the Fundamental Research Funds for the Central Universities of China No. YJ202343. Wyman was supported by NSF grant DMS-2204397. Wyman extends a thanks to Richard Lange for his very patient guidance through setting up a GitHub repository.
@@ 2. Closed walks of bounded length determine if two vertices are cospectral
We note that the number $n-1$ of counts is sharp since the adjacency matrix $A$ of a simple path with $n$ vertices for every odd $n=2\ell+1$ (i.e., $A_{i,j}=\delta_{|i-j|=1}$) as well as the adjacency matrix $A$ of a simple path with $n$ vertices and a self-loop attached to the last vertex for every even $n=2\ell$ (i.e., $A_{i,j}=\delta_{|i-j|=1}+\delta_{i=j=n}$) satisfies
@@ 3. Minimal examples
The code is written in the Julia language (see [BEKS17, FBS+21]) and contains a number of tools:
@@
    A number of other convenience functions, such as those that determine if a graph is walk-regular or if there exist non-similar cospectral pairs.
@@ 3.3. Minimal walk-regular, non-vertex-transitive graphs
There are no walk-regular, non-vertex-transitive graphs, which have less than 12 vertices. There are exactly four minimal examples with twelve vertices, one each of degree 4, 5, 6, and 7. Again, these minimal examples are connected, and so can be put into edge-complement pairs; the examples of degrees $4$ and $7$ are edge complements, and so are the examples of $5$ and $6$. The degree $4$ and $5$ examples are depicted in Figure 3. Furthermore, even though the degree 4 example is not planar, it can be embedded into a torus, as shown in Figure 4.
@@
Figure 4. A walk-regular non-vertex-transitive toroidal graph.
@@ 4. Walk-regular planar graphs
4.1. Auxiliary lemmas
Here we set up the tools we will need in order to classify 3-connected walk-regular planar graphs and prove Theorem 1.5.
The first tool shall be used to obtain the number of short closed walks through a vertex from the number of cycles through it. Let $G$ be a $d$-regular graph with $d\geq 3$. A closed walk is said to be contractible if it traces no cycles. For a vertex $v$ and $m\geq 3$, we let
$\mathcal{C}_{m}(v)=\#\{\text{$m$-cycles through $v$}\}.$
As every closed walk of length 3 traces a triangle in one of two directions,
$(A^{3})_{v,v}=2\mathcal{C}_{3}(v).$
Starting at $v$ there are exactly $d(d-1)+d^{2}$ contractible closed walks with length 4. Adding in the ones tracing 4-cycles, we have
$(A^{4})_{v,v}=2d^{2}-d+2\mathcal{C}_{4}(v).$
Now we assume that $\mathcal{C}_{3}(v_{i})\equiv\mathcal{C}_{3}$ for all vertices $v_{i}$. Note that every closed walk $\overline{va_{1}a_{2}a_{3}a_{4}v}$ traces either a 5-cycle or an edge attached to a triangle; among the latter ones, there are exactly $2d\mathcal{C}_{3}$ with $v=a_{2}$, $2d\mathcal{C}_{3}$ with $v=a_{3}$, $2(d-2)\mathcal{C}_{3}$ with $a_{1}=a_{4}$ and $v\notin\{a_{2},a_{3}\}$, $2(d-2)\mathcal{C}_{3}$ with $a_{1}=a_{3}$ and $a_{2}\notin\{v,a_{3}\}$, $2(d-2)\mathcal{C}_{3}$ with $a_{2}=a_{4}$ and $a_{3}\notin\{v,a_{1}\}$, and $2\mathcal{C}_{3}$ with $a_{1}=a_{3}$ and $a_{2}=a_{4}$. In total we have
$(A^{5})_{v,v}=10(d-1)\mathcal{C}_{3}+2\mathcal{C}_{5}(v).$
Furthermore, assume that $\mathcal{C}_{4}(v_{i})\equiv\mathcal{C}_{4}$ for all vertices $v_{i}$. If $G$ is triangle-free (i.e., $\mathcal{C}_{3}=0$), similarly, we count the closed walks starting at $v$ with length 6: $d^{3}+2d^{2}(d-1)+2d(d-1)^{2}$ of them are contractible, $2\cdot 2d\mathcal{C}_{4}+4\cdot 2(d-2)\mathcal{C}_{4}+2\cdot 2\mathcal{C}_{4}$ of them trace some 4-cycles with attached edges, and the remaining $2\mathcal{C}_{6}(v)$ trace some 6-cycles. Therefore,
$(A^{6})_{v,v}=5d^{3}-6d^{2}+2d+12(d-1)\mathcal{C}_{4}+2\mathcal{C}_{6}(v).$
If $\mathcal{C}_{3}=1$, compared to the triangle-free case, there are two fewer contractible closed walks of length 6 but four more such walks that trace the triangle through $v$ twice. In this case, it follows that
$(A^{6})_{v,v}=5d^{3}-6d^{2}+2d+12(d-1)\mathcal{C}_{4}+2\mathcal{C}_{6}(v)+2.$
A direct consequence of the above formulas is that
Lemma 4.1.
Let $G$ be a walk-regular graph. Then
  1. (1)
    $\mathcal{C}_{3}(v)\equiv\mathcal{C}_{3},\,\mathcal{C}_{4}(v)\equiv\mathcal{C}_{4}$, and $\mathcal{C}_{5}(v)\equiv\mathcal{C}_{5}$ for all vertices $v_{i}$;
  2. (2)
    if $\mathcal{C}_{3}\leq 1$, $\mathcal{C}_{6}(v)$ is also constant over all vertices $v_{i}$.
@@ 4.1. Auxiliary lemmas
Using Euler’s identity $n+f-e=2$ and $e=nd/2$, we get
(4.1) $\frac{1}{n}\sum_{v}p(v)=\frac{d-2}{2}+\frac{2}{n}.$
In order to prove Theorem 1.5, we shall prove the following key lemma.
@@
For any 3-connected walk-regular planar graph with degree $d\geq 3$, the modified Schläfli symbol is constant over all vertices, that is
@@
However, establishing (4.2) succinctly appears challenging. Instead, we will justify (4.2) in each individual case.
We remark that, once we prove Lemma 4.2, it is routine to check that Theorem 1.5 holds in each case. Indeed, the numbers of vertices, edges, and faces are determined by Lemma 4.2 and Euler’s identity. The graph can then be reconstructed in a deterministic fashion, starting with an arbitrary polygon as the outer face. We illustrate this fact for the most complicated case when $r(v)\equiv(3,4,5,4)$ in Figure 5. The only exception is the case when $r(v)\equiv(3,4,4,4)$, for which one runs into two possible reconstructions, the polyhedral graphs of the $(3,4,4,4)$-solid and twisted $(3,4,4,4)$-solid. However, the twisted one is not walk-regular. See Case (13) in the proof.
@@
We fix a vertex $v_{1}$ such that $p(v_{1})=\max_{v}p(v)$. Let
$r_{1}=r_{1}(v_{1})\quad\text{and}\quad k=\#\{1\leq i\leq d:r_{i}(v_{1})=r_{1}(v_{1})\}.$
In view of (4.1), to prove Lemma 4.2, we only need to discuss the fifteen cases enumerated below. For the convenience of the reader, we have also included an exhaustive list of the possible walk-regular planar graphs belonging to each case. From now on and throughout the paper, for the purpose of simplicity, we shall not distinguish between a solid and its polyhedral graph.
@@
Table 1. 15 possible cases.
To effectively use Lemma 4.1 in discussions, we need another tool to convert the number of short cycles into the one of faces. We label an $m$-cycle $C=\overline{a_{1}a_{2}\cdots a_{m}a_{1}}$ by its inner degree sequence
$d_{*}(C)=(d_{*}(a_{1}),d_{*}(a_{2}),\cdots,d_{*}(a_{m})),$
where $d_{*}(a_{i})$ is the number of incident edges of $a_{i}$ inside $C$; here an object (vertex, edge, cycle, face, etc) is said to be “inside $C$” if it lies in the region bounded by $C$ in the planar embedding. Again, $d_{*}(C)$ is uniquely defined, once we choose $a_{1}$ and the direction to minimize the sequence lexicographically. Clearly, $C$ is a face if and only if $d_{*}(C)=(0,0,\cdots,0)$ or $(d-2,d-2,\cdots,d-2)$ (the latter means that $C$ is the unique outer face of the embedding). For $m=3$, we have
Lemma 4.3.
In a 3-connected walk-regular planar graph, every triangle is a face.
Note that a cycle $C$ in $G$ with length $\geq 4$ is not necessarily a face. The typical situation is that $C$ has a chord, an edge not in $C$ whose endpoints lie in $C$. For instance, the (3,3,3,3,4)-solid contains a 4-cycle with one chord, a 5-cycle with one chord and a 5-cycle with two chords; none of them are faces. We also observe that the 4-cycle encircling four triangles in the octahedron, the 5-cycle encircling five triangles in the icosahedron, and the 5-cycle encircling three triangles and a quadrangular face in the 4-gonal antiprism are not faces, even though all of them are chordless. Indeed, by excluding these cases, we have the next two lemmas for 4- and 5-cycles.
Lemma 4.4.
In a 3-connected walk-regular planar graph which is not isomorphic to the octahedron, every chordless 4-cycle is a face.
Lemma 4.5.
In a 3-connected walk-regular planar graph which is not isomorphic to the 4-gonal antiprism or the icosahedron, every chordless 5-cycle is a face.
Next, we will provide the proofs of Lemmas 4.34.5 separately in the cases $d=3,4,5$, and conclude the proof of Lemma 4.2, and thus Theorem 1.5 for each case.
4.2. The cubic cases
Assume that $G$ is cubic (i.e., $3$-regular). When a chordless $m$-cycle $C$ in $G$ is not a face, the $3$-connectedness yields that $\geq 3$ vertices on $C$ have incident edges inside $C$, and $\geq 3$ of them have incident edges outside $C$, hence $m\geq 6$. This justifies Lemmas 4.34.5 for all cubic cases. Consequently, for any vertex $v$, we have
$r_{1}(v)\equiv r_{1}\quad\text{and}\quad k(v):=\#\{1\leq i\leq 3:r_{i}(v)=r_{1}\}\equiv k.$
Among the cases (1)–(9) in Table 1, (2) and (8) are simply impossible to realize. Also, one quickly sees the symbols $r(v)$ of the graphs realizing cases (1), (4), and (7) must be constant, where the corresponding determined solids are the tetrahedron, the hexahedron, and the dodecahedron.
In case (5) ($d=3,r_{1}=4,k=2$), the face $f_{3}(v_{1})$ surrounded by $m=r_{3}(v_{1})\geq 5$ quadrangular faces determines the unique pattern up to isomorphism. Therefore, $r(v)\equiv(4,4,m)$ and the graph is isomorphic to the $m$-gonal prism.
For the other three cases (with $d=3,k=1$), we need to consider the second shortest cycles in $G$, that is, the cycles with length
$\ell=\min\{m>r_{1}:\mathcal{C}_{m}(v)>0\text{ for some vertex }v\}.$
It is clear that $\ell\leq r_{2}(v_{1})$. We will claim in each individual case that
Lemma 4.6.
In a 3-connected walk-regular planar graph with $d=3$ and $k=1$, every $\ell$-cycle is a face.
@@ 4.2. The cubic cases
($d=3,r_{1}=3,k=1$) Note that each $\ell$-cycle is chordless, and it should be adjacent to exactly $\ell/2$ triangles, and thus $\ell$ must be even. Suppose that an $\ell$-cycle $C$ is not a face. Then by the 3-connectedness, $C$ has $\geq 3$ neighbors inside $C$ and $\geq 3$ neighbors outside $C$. Since $C$ is chordless and each vertex must be adjacent to a triangle, we must have $\ell\geq 12$. However, in this situation one has $r_{2}(v_{1})\geq 12$ and
$p(v_{1})\leq\frac{1}{3}+\frac{1}{12}+\frac{1}{12}=\frac{1}{2},$
which violates (4.1). This proves Lemma 4.6 for $r_{1}=3$, and the possible $\ell=4,6,8$ or $10$.
Assume $\ell=4$ or $6$. It follows from Lemma 4.1 that $\mathcal{C}_{4}(v)$ or $\mathcal{C}_{6}(v)$ is constant for all vertices $v$. For $\ell=8$, we can count the number of $8$-cycles though a vertex in its 4-neighborhood to obtain
$(A^{8})_{v,v}=591+2\mathcal{C}_{8}(v).$
For $\ell=10$, a similar counting in the 5-neighborhood of the vertex $v$ provides
$(A^{10})_{v,v}=4223+2\mathcal{C}_{10}(v).$
Hence whatever $\ell$ is, we have $\mathcal{C}_{\ell}(v)\equiv\mathcal{C}_{\ell}$. Since 3 is odd, drawing a triangular face on which each vertex is adjacent to exactly one face with length $\ell$ is impossible. The only other possibility is $\mathcal{C}_{\ell}=2$, and it follows that $r(v)\equiv(3,\ell,\ell)$. The corresponding solids are the triangular prism, the $(3,6,6)$-solid, the $(3,8,8)$-solid and the $(3,10,10)$-solid respectively.
@@
($d=3,r_{1}=4,k=1$) We first claim that every cycle has even length. If this was false, we let $\ell^{\prime}=\min\{\text{odd }m:\mathcal{C}_{m}(v)>0\text{ for some vertex }v\}$, and let $C$ be a minimal $\ell^{\prime}$-cycle, in the sense that no $\ell^{\prime}$-cycles are properly contained in the closed region bounded by $C$. Suppose that a vertex $b_{1}$ is inside $C$. The walk-regularity implies that, starting at $b_{1}$ there is at least one closed walk with length $\ell^{\prime}$, denoted by $C^{\prime}=\overline{b_{1}b_{2}\cdots b_{\ell^{\prime}}b_{1}}$. By the minimality of $\ell^{\prime}$ and $C$, $C^{\prime}$ should be a cycle, and some of vertices on $C^{\prime}$ lie outside $C$. Let $b_{i},b_{j}$ be the first and the last intersection vertices of $C$ and $C^{\prime}$ respectively, that is, $i=\min\{m:b_{m}~{}\text{lies on}~{}C\}$ and $j=\max\{m:b_{m}~{}\text{lies on}~{}C\}$. Such $b_{i}$ and $b_{j}$ divide $C$ into two paths $\mathcal{P}_{1},\mathcal{P}_{2}$ with length $\ell_{1},\ell_{2}$, and since $\ell_{1}+\ell_{2}=\ell^{\prime}$ is odd, one of $\ell_{1},\ell_{2}$ (say $\ell_{1}$) has the same parity as $\ell^{\prime}+i-j$. As the union of two paths $\overline{b_{i}b_{i-1}\cdots b_{1}b_{\ell^{\prime}}b_{\ell^{\prime}-1}\cdots b_{j}}$ and $\mathcal{P}_{2}$ form a cycle with odd length $\ell^{\prime}+i-j+\ell_{2}$ inside the closed region bounded by $C$, it follows again from the minimality of $\ell^{\prime}$ and $C$ that $\ell^{\prime}+i-j+\ell_{2}>\ell^{\prime}$. But now, one has the union of two paths $\overline{b_{i}b_{i+1}\cdots b_{j}}$ and $\mathcal{P}_{1}$ as a closed walk with odd length $j-i+\ell_{1}<\ell^{\prime}$, which still violates the minimality of $\ell^{\prime}$. Therefore, $C$ has no inner neighbors, and it should be an inner face in the planar embedding; however, this is also impossible, since at least one adjacent 4-cycle of $C$ shares three vertices with $C$, of which the middle one must have an inner incident edge contradicting Lemma 4.4. The claim is justified.
Using (4.1), we obtain $\ell\leq r_{2}(v_{1})<8$. By the claim, the only possibility is $\ell=r_{2}(v_{1})=6$. For a 6-cycle $C$, it has either two or three adjacent $4$-cycles. In the former situation, $C$ has only two inner neighbors and two outer neighbors, which violates 3-connectedness. For the latter one, to be 3-connected all of the three adjacent $4$-cycles must lie outside $C$ whenever $C$ is not the outer face. This proves that every $6$-cycle is a face, and justifies Lemma 4.6 for $r_{1}=4$.
In view of Lemma 4.1, $\mathcal{C}_{6}(v)\equiv\mathcal{C}_{6}=1$ or $2$. For $\mathcal{C}_{6}=2$, we have $r(v)\equiv(4,6,6)$, and the graph is isomorphic to the $(4,6,6)$-solid.
Now we consider the sub-case $\mathcal{C}_{6}=1$. It follows from (4.1) and the claim that $r_{3}(v_{1})=8$ or $10$. For a vertex $v$ with $r(v)=(4,6,m)$, by counting in the 4- and 5-neighborhood of $v$ we obtain
$(A^{8})_{v,v}=809+\begin{cases}\,2\quad\text{if }m=r_{3}(v_{1})=8,\\ \,0\quad\text{if }m>r_{3}(v_{1})=8,\end{cases}$
and
$(A^{10})_{v,v}=6063+\begin{cases}\,2\quad\text{if }m=r_{3}(v_{1})=10,\\ \,0\quad\text{if }m>r_{3}(v_{1})=10.\end{cases}$
Hence, $r(v)\equiv(4,6,8)$ or $\equiv(4,6,10)$, which corresponds to the $(4,6,8)$-solid or the $(4,6,10)$-solid respectively.